OpenStudy (anonymous):

Solve y = 3x – 12 by the substitution method. 2x – 3y = -13 Answer (7, 9) (4, 3) (1, 5) (0, 8)

OpenStudy (anonymous):

substitute 3x-12 for y so 2x-3(3x-12)=-13 Can you get the answer from this equation?

TheSmartOne (thesmartone):

So just substitute 2x-3(3x-12)=-13 2x-9x+36=-13 -7x+36=-13 -7x=-49 x=-7

TheSmartOne (thesmartone):

Oops x=7 not -7

TheSmartOne (thesmartone):

And since A is the only one with 7 as the x value, A is the answer